SCALES OF MEASUREMENT VERSION 1.8 William Nielsen Brandt — niel@astro.psu.edu

3 × 10^{−6} arcseconds Angular size of a 10^{8} M_{⊙} Schwarzschild black hole in M31
0.0002  arcseconds Typical VLBA resolution at 43 GHz
0.0005  arcseconds Person on the Moon
0.001  arcseconds Typical VLBI resolution
0.001  arcseconds Typical gravitational microlensing image angular separation
0.0015  arcseconds Hipparcos satellite astrometric positional accuracy
0.04  arcseconds Hubble Space Telescope FOC optical angular resolution
0.1  arcseconds VLA largest configuration resolution at 23 GHz
0.1  arcseconds Hubble Space Telescope WFC optical angular resolution
0.15  arcseconds 1 meter optical reflector yellow light diffraction limit
0.5  arcseconds WHT MARTINI guide star system resolution
0.6  arcseconds Rough optical telescope atmospheric seeing limit
0.8  arcseconds Angular diameter of Europa
1  arcseconds AXAF X-ray satellite angular resolution goal
1.32  arcseconds Yearly proper motion of Sirius
1.75  arcseconds Gravitational deflection of a light ray which grazes the Sun
2.9  arcseconds Angular diameter of a M31 sized galaxy at z = 0.5 in optical light
5  arcseconds ROSAT X-ray satellite HRI angular resolution
8  arcseconds Angular separation of the gravitationally lensed quasar 0957+561
25  arcseconds ROSAT X-ray satellite PSPC angular resolution
36  arcseconds Angular diameter of Jupiter
120  arcseconds Human eye resolution
120  arcseconds Typical optical telescope field of view
150  arcseconds Rough Crab Nebula radius
210  arcseconds Rough Cir X-1 radio nebula radius
400  arcseconds Cluster imprint on the cosmic microwave background radiation
480  arcseconds OVRO 5.5 meter 32 GHz telescope resolution
700  arcseconds Jodrell bank 250 foot 1 420.4 MHz telescope resolution
1 200  arcseconds ROSAT X-ray satellite HRI field of view radius
1 500  arcseconds ASCA X-ray satellite GIS field of view radius
1 865  arcseconds Angular diameter of the Moon at its mean distance
1 800  arcseconds 100 Mpc structure imprint on the cosmic microwave background radiation
1 922  arcseconds Angular diameter of the Sun at its mean distance
2 700  arcseconds Largest dimension of W50 (ARAA 22, 524)
3 100  arcseconds Maximum angular scale of causal connection on the cosmic microwave background radiation
3 600  arcseconds ROSAT X-ray satellite PSPC field of view radius
10 800  arcseconds Angular diameter of the LMC
14 400  arcseconds Angular diameter of M31 in optical light
19 800  arcseconds Palomar Optical Sky Survey plate field of view
25 200  arcseconds COBE DMR resolution limit
5 × 10^{−44} seconds Planck-Wheeler time = (\frac{Gh̄}{c^{5}})^{\frac{1}{2}}
4 × 10^{−24} seconds Typical lifetime of strong interaction resonance = \frac{h}{m_{p}c^{2}}
8 × 10^{−21} seconds Electron light crossing time = \frac{h}{m_{e}c^{2}}
1 × 10^{−13} seconds Typical period of vibration of an atom in a solid
1 × 10^{−13} seconds Typical X-ray line electric dipole radiative transition time
1.6 × 10^{−9} seconds Typical hydrogen 2p→1s radiative transition time (electric dipole one photon process)
8 × 10^{−4}  seconds Mass shedding minimum spin period for a neutron star
1.6 × 10^{−3} seconds Spin period of PSR 1957+20
0.12  seconds Typical hydrogen 2s→1s radiative transition time (strictly forbidden two photon process)

10  seconds Median duration of a classical γ-ray burst
887  seconds Mean life of a neutron in free space
2 000  seconds Sun dynamic time scale
8.6 × 10^{4}  seconds Earth rotation time
3.2 × 10^{7}  seconds Earth orbit time around the Sun
1.6 × 10^{9}  seconds Typical time between Milky Way supernovae
1.9 × 10^{11} seconds Carbon-14 half-life
3 × 10^{12}  seconds Rough lifetime of a supernova remnant
1.5 × 10^{13} seconds Typical HMXB evolution time
3 × 10^{13}  seconds Rough time for evolution of a biological species
1.6 × 10^{14} seconds Main sequence lifetime for a 30 M_{⊙} star
3 × 10^{14}  seconds Hydrogen 21 cm spin flip time
3 × 10^{14}  seconds Rough Lyapunov time of the solar system
6.3 × 10^{14} seconds Sun thermal time scale
1.3 × 10^{15} seconds E-folding time for a black hole accreting at the Eddington rate with an efficiency of 0.1
1.9 × 10^{15} seconds Main sequence lifetime for a 5 M_{⊙} star
2 × 10^{15}  seconds Timescale for Los Angeles to pass San Francisco via continental drift
2.4 × 10^{15} seconds Typical LMXB evolution time
7.3 × 10^{15} seconds Orbit time for sun around galaxy center
1.1 × 10^{16} seconds PSR 1913+16 orbital gravitational radiation coalescence timescale
2 × 10^{16}  seconds Rough supernova biological extinction time
6 × 10^{16}  seconds Minimum age of PSR J0437–4715 (ApJ 411, L85)
6 × 10^{16}  seconds Time for galaxy to cross a cluster
7 × 10^{16}  seconds Radial period of the orbit of the Magellanic Clouds (MNRAS 198, 718)
1.1 × 10^{17} seconds Primeval slime to man time
1.5 × 10^{17} seconds Age of Earth and Sun
1.5 × 10^{17} seconds Uranium-238 half-life
2.7 × 10^{17} seconds Look back time to z = 1
3 × 10^{17}  seconds Main sequence lifetime for a 1 M_{⊙} star
3.3 × 10^{17} seconds Look back time to z = 2
3.3 × 10^{17} seconds Sun nuclear time scale
3.7 × 10^{17} seconds Look back time to z = 4
3.8 × 10^{17} seconds Rough age of the Milky Way
3.8 × 10^{17} seconds Look back time to the z = 4.897 quasar PC 1247+3406
4 × 10^{17}  seconds Rough age of 47 Tucanae (an old globular cluster)
4.1 × 10^{17} seconds Age of the universe = \frac{2}{3H_{0}}
2.5 × 10^{18} seconds Globular cluster evaporation time
3 × 10^{25}  seconds Galaxy dynamical relaxation timescale
3 × 10^{27}  seconds Earth/Sun orbital gravitational radiation coalescence timescale
1 × 10^{39}  seconds Lower limit on the proton lifetime
4.7 × 10^{73} seconds 1 M_{⊙} black hole Hawking evaporation time
1.6 × 10^{−35}  meters Planck-Wheeler length = (\frac{Gh̄}{c^{3}})^{\frac{1}{2}}
2 × 10^{−35}  meters Rough postulated superstring size
1 × 10^{−24}  meters Effective 1 MeV neutrino radius
1.6 × 10^{−22}  meters Radius of a 10 M_{⊙} star squeezed down to the Planck-Wheeler density
1.5 × 10^{−18}  meters Classical proton radius = \frac{e^{2}}{4πϵ_{0}m_{p}c^{2}}
1.6 × 10^{−17}  meters W^{±} Compton wavelength = \frac{h}{m_{W^{±}} c} = rough weak force length
4 × 10^{−17}  meters LIGO 4 km gravity-wave detector needed sensitivity
1.3 × 10^{−15}  meters Proton Compton wavelength = \frac{h}{m_{p}c} = rough proton fuzziness length
1.44 × 10^{−15}A^{\frac{1}{3}} meters Nuclear radius
2.8 × 10^{−15}  meters Classical electron radius = \frac{e^{2}}{4πϵ_{0}m_{e}c^{2}}

8.8 × 10^{−15}  meters π-meson Compton wavelength = \frac{h}{m_{π}c} = attractive strong force length
1.2 × 10^{−12}  meters 1 MeV γ-ray wavelength
2.4 × 10^{−12}  meters Electron Compton wavelength = \frac{h}{m_{e}c} = rough electron fuzziness length
5.3 × 10^{−11}  meters Bohr radius = \frac{h^{2}ϵ_{0}}{πm_{e}e^{2}}
2.6 × 10^{−10}  meters Copper atom spacing in solid copper
3.5 × 10^{−10}  meters H_{2}O molecular diameter
4 × 10^{−10}  meters ROSAT X-ray satellite mirror rms surface error
1.2 × 10^{−9}  meters 1 keV X-ray wavelength
3 × 10^{−9}  meters Typical mean nucleon spacing during primordial nucleosynthesis
3.4 × 10^{−9}  meters DNA double helix turn length
6 × 10^{−8}  meters Typical rms surface error of a Keck telescope mirror
7.0 × 10^{−8}  meters Molecular mean free path in the atmosphere
1 × 10^{−7}  meters Typical size of a virus
3 × 10^{−7}  meters Interstellar dust grain size
5 × 10^{−7}  meters Optical photon wavelength
4 × 10^{−6}  meters Typical size of a cell
2 × 10^{−4}  meters Small dust particle size
0.03  meters Lunar laser ranging accuracy (Science 265, 482)
0.068  meters Unraveled human DNA strand length
1.8  meters Man
30  meters Blue Whale
3 700  meters Mean ocean depth
4 000  meters Error in our knowledge of the distance to Jupiter
5 500  meters Rough radius of Halley’s comet
8 847  meters Height of Mount Everest
10 000  meters Neutron star radius
10 000  meters Typical asteroid radius
10 000  meters Typical comet radius
11 032  meters Depth of the Mariana Trench
12 000  meters Height of troposphere
12 000  meters Typical airliner cruising altitute
30 000  meters Typical thickness of the Earth’s crust
5.5 × 10^{5}  meters Height of the ASCA X-ray satellite orbit
9 × 10^{5}  meters Height of the COBE satellite orbit
3.2 × 10^{6}  meters Length of the Great Wall of China
6.3 × 10^{6}  meters Radius of the Earth
4.2 × 10^{7}  meters Geostationary satellite orbit height
7.1 × 10^{7}  meters Radius of Jupiter
9 × 10^{7}  meters Distance to the Earth’s solar wind bow shock
3.8 × 10^{8}  meters Distance to the Moon
7.0 × 10^{8}  meters Radius of the Sun
1 × 10^{9}  meters Typical X-ray binary accretion disk circularization radius
1.7 × 10^{9}  meters 5 M_{⊙} star main sequence radius
7 × 10^{9}  meters Typical standoff radius of Jupiter’s solar wind bow shock
1.47 × 10^{11}  meters Event horizon radius for a 10^{8} M_{⊙} maximally rotating Kerr black hole
1.50 × 10^{11}  meters Earth/Sun mean distance
1.5 × 10^{11}  meters Radius of the red giant Mira at minimum light
3 × 10^{11}  meters Radius of the red giant Mira at maximum light
5 × 10^{11}  meters Radius of the 20 M_{⊙} red supergiant Betelgeuse at maximum light
5.91 × 10^{12}  meters Pluto/Sun mean distance
1.5 × 10^{13}  meters Expected distance to the solar wind termination shock
2 × 10^{14}  meters Rough stellar separation in the central parsec of the Milky Way
4 × 10^{14}  meters Rough stellar separation in a globular cluster

4 × 10^{14}  meters Seyfert galaxy characteristic broad line region radius
2 × 10^{15}  meters Rough Oort Cloud/Sun distance
9.46 × 10^{15}  meters 1 light-year
3.08 × 10^{16}  meters 1 parsec
4 × 10^{16}  meters Nearest nonsolar star to Earth
4.5 × 10^{16}  meters Rough Crab Nebula radius
1.6 × 10^{17}  meters Stromgren sphere radius for an O5 star
3 × 10^{17}  meters Rough Cir X-1 radio nebula radius
3 × 10^{17}  meters Rough supernova biological extinction distance (PASP 106, 689)
5 × 10^{17}  meters Typical interstellar medium cloud size
1.4 × 10^{18}  meters Hyades open cluster distance
1.5 × 10^{18}  meters Trigonometric parallax distance determination limit
1.5 × 10^{18}  meters Typical globular cluster radius
2 × 10^{18}  meters Rough radius of the local interstellar hot gas bubble
2 × 10^{18}  meters Largest dimension of W50 (ARAA 22, 524)
2.6 × 10^{18}  meters Rough distance to PSR J0108 − 1431
5 × 10^{18}  meters Scale height of the Milky Way interstellar medium
5 × 10^{18}  meters Seyfert galaxy characteristic narrow line region radius
5.2 × 10^{18}  meters Distance to the 20 M_{⊙} red supergiant Betelgeuse (will go supernova within 10 000 years)
2 × 10^{19}  meters Characteristic height of the Milky Way main disk
6 × 10^{19}  meters Typical dwarf galaxy radius
6.6 × 10^{19}  meters Distance to the Crab Nebula
2 × 10^{20}  meters Estimated distance to Cir X-1
2.4 × 10^{20}  meters Distance from Sun to galactic center
1.5 × 10^{21}  meters Distance to the LMC
2 × 10^{21}  meters Rough Milky Way dark matter halo radius
6 × 10^{21}  meters Typical Ly-α galaxy neutral hydrogen radius
1 × 10^{22}  meters Typical active galaxy jet length
1 × 10^{22}  meters Length of the “superantennae” of IRAS 19254–7245
1.9 × 10^{22}  meters Distance to M31
3 × 10^{22}  meters Radius of the core of the Virgo cluster of galaxies
4 × 10^{22}  meters Rough Local Group radius
9.2 × 10^{22}  meters Cepheid variable distance determination limit (via ground observations)
1.2 × 10^{23}  meters Distance to the Circinus galaxy
5 × 10^{23}  meters Rough boundary between the nonlinear and linear gravitational collapse regimes
7 × 10^{23}  meters Distance to the z = 0.0036 Seyfert 2 NGC 1068
7 × 10^{23}  meters Distance to the center of the Virgo cluster of galaxies
9 × 10^{23}  meters Characteristic maximum travel length of a yocto-eV cosmic ray due to pion

production off the microwave background

2.3 × 10^{24}  meters Shapley supercluster radius
3 × 10^{24}  meters Typical length scale probed by the Harvard/Smithsonian CfA redshift survey
4.1 × 10^{24}  meters Luminosity distance to the z = 0.022 Coma cluster of galaxies
1.5 × 10^{25}  meters Typical length scale probed by the Lick survey
3.0 × 10^{25}  meters Luminosity distance to the z = 0.158 quasar 3C273
3 × 10^{25}  meters Typical length scale probed by the 4C radio galaxy survey
3 × 10^{25}  meters Schwarzschild radius of a singularity with the mass of a critical density universe
5.5 × 10^{26}  meters Luminosity distance to the z = 2.286 ultraluminous IRAS galaxy IRAS F10214+4724
1.3 × 10^{27}  meters Luminosity distance to the z = 4.897 quasar PC 1247+3406
1 × 10^{−9}  m s^{−1} Sea floor spreading rate
1.6 × 10^{−9} m s^{−1} Average slip rate of the San Andreas fault
1 × 10^{−8}  m s^{−1} Typical rainfall rate in a semi-arid climate
2 × 10^{−8}  m s^{−1} Grass growth rate
3 × 10^{−6}  m s^{−1} Typical glacial advance rate

1 × 10^{−3}  m s^{−1} Equivalent radial velocity resolution of pulsar pulse arrival time analysis
1.3  m s^{−1} Human walking speed
3  m s^{−1} Radial velocity accuracy of high precision Doppler spectroscopy
13  m s^{−1} Speed of the reflex motion induced on the Sun by Jupiter
25  m s^{−1} Car speed
60  m s^{−1} Radial velocity semiamplitude of 51 Pegasi due to its planet
100  m s^{−1} Typical speed of an electric pulse in the nervous system
330  m s^{−1} Sound speed in air
480  m s^{−1} Earth’s atmosphere molecular rms velocity
600  m s^{−1} Fighter jet speed
2 380  m s^{−1} Escape velocity from Moon’s surface
10 000  m s^{−1} Typical longitudinal seismic wave velocity in the Earth’s mantle
11 000  m s^{−1} Escape velocity from the Earth’s surface
20 000  m s^{−1} Globular cluster stellar velocity dispersion
29 000  m s^{−1} Earth’s motion around the Sun
40 000  m s^{−1} Globular cluster stellar escape velocity
1 × 10^{5}  m s^{−1} Typical Galactic pulsar vertical velocity component
1 × 10^{5}  m s^{−1} Average speed of the initial stroke of a lightning flash
2.2 × 10^{5}  m s^{−1} Rotational velocity of the Sun around the Milky Way’s center
3 × 10^{5}  m s^{−1} Orbital speed of PSR 1913+16
3 × 10^{5}  m s^{−1} Rough velocity of Geminga’s proper motion
3.65 × 10^{5} m s^{−1} Motion of the solar system barycenter relative to the cosmic microwave background
4.1 × 10^{5}  m s^{−1} Orbital speed of Cen X-3
6.2 × 10^{5}  m s^{−1} Escape velocity from the Sun’s surface
6.2 × 10^{5}  m s^{−1} Escape velocity from the Milky Way for objects in the solar neighborhood (ARAA 29, 429)
6.22 × 10^{5} m s^{−1} Motion of the Local Group relative to the cosmic microwave background
8 × 10^{5}  m s^{−1} Typical galaxy cluster galaxy velocity dispersion
2 × 10^{6}  m s^{−1} Speed of n = 1 hydrogen electron = Zαc
5 × 10^{6}  m s^{−1} Young (months old) supernova ejecta
1.2 × 10^{7}  m s^{−1} Velocity of the wind streaming out from H1413+117 “Cloverleaf” (Nature 371, 559)
7.8 × 10^{7}  m s^{−1} SS433 jet speed
1.4 × 10^{8}  m s^{−1} Keplerian orbital velocity at the surface of a neutron star
2 × 10^{8}  m s^{−1} Escape velocity from neutron star surface
2.8 × 10^{8}  m s^{−1} Ejecta from GRS 1915+105 (Nature 371, 46)
2.998 × 10^{8} m s^{−1} Light in a vacuum
3.7 × 10^{8}  m s^{−1} Apparent superluminal motion of the ejecta from GRS 1915+105 (Nature 371, 46)
6 × 10^{9}  m s^{−1} Apparent superluminal motion of the jet from the z = 2.16 quasar 0836+710
2.7 × 10^{10} m s^{−1} Apparent superluminal motion of the jet from the z = 0.940 BL Lac AO 0235+164 (may be lensed)
4.2 × 10^{−36} kilograms Mass equivalent of a green light photon
9.1 × 10^{−36} kilograms Electron antineutrino upper mass limit
4.8 × 10^{−31} kilograms Muon neutrino upper mass limit
9.11 × 10^{−31} kilograms Electron mass
5.5 × 10^{−29} kilograms Tau neutrino upper mass limit
1.67 × 10^{−27} kilograms Proton mass
9 × 10^{−27}  kilograms Bottom quark mass
4.8 × 10^{−26} kilograms Mean mass of atmosphere molecule
1.4 × 10^{−25} kilograms W^{±} mass (80.22 ± 0.22 GeV/c^{2})
1.6 × 10^{−25} kilograms Z^{0} mass (91.187 ± 0.007 GeV/c^{2})
2 × 10^{−25}  kilograms Favored Higgs boson mass
3.1 × 10^{−25} kilograms Top quark mass (176 ± 13 GeV/c^{2})
4 × 10^{−25}  kilograms DNA nucleotide
1 × 10^{−22}  kilograms Typical protein molecule mass
5 × 10^{−21}  kilograms E. Coli ribosome

1 × 10^{−16}  kilograms Interstellar dust grain mass
8 × 10^{−15}  kilograms Rough mass of a human DNA molecule
7 × 10^{−13}  kilograms Typical mass of a cell
2.2 × 10^{−8}  kilograms Planck-Wheeler mass = (\frac{h̄c}{G})^{\frac{1}{2}}
1 × 10^{−5}  kilograms Typical mosquito mass
0.02  kilograms Typical goldfish mass
70  kilograms Typical human mass
70  kilograms Lower limit to the allowed mass for a Sumo wrestler
100  kilograms Meteorite mass before entry into Earth’s atmosphere
420  kilograms ASCA X-ray satellite mass
900  kilograms Compton Gamma-Ray Observatory satellite mass
1 000  kilograms Car
2 200  kilograms ISO infrared satellite mass at launch
10 000  kilograms Tyrannosaurus Rex
5 × 10^{11}  kilograms Initial mass of a primordial black hole with evaporation time equal to the universe’s age
1 × 10^{13}  kilograms Typical comet mass
3 × 10^{14}  kilograms Typical mountain mass
1.1 × 10^{16}  kilograms Superterranean biomass of Earth (ocean organisms are included)
5.3 × 10^{18}  kilograms Total mass of Earth’s atmosphere
3 × 10^{19}  kilograms Typical asteroid mass
1.4 × 10^{21}  kilograms Total mass of Earth’s oceans
4.8 × 10^{22}  kilograms Mass of Europa
7.3 × 10^{22}  kilograms Mass of the Moon
2 × 10^{24}  kilograms Rough Oort cloud mass
5.98 × 10^{24}  kilograms Mass of the Earth
1.9 × 10^{27}  kilograms Mass of Jupiter
1.6 × 10^{29}  kilograms Minimum mass to fusion burn hydrogen
4 × 10^{29}  kilograms Rough mass of MACHO and EROS microlensing objects
1.99 × 10^{30}  kilograms Mass of the Sun
2.8 × 10^{30}  kilograms Chandrasekhar mass (maximum mass for a white dwarf)
6.0 × 10^{30}  kilograms Oppenheimer-Volkoff mass (maximum mass for a neutron star)
1.2 × 10^{31}  kilograms Minimum mass of the unseen object in GS2023+338/V404 Cygni
4 × 10^{31}  kilograms Rough stellar mass above which the evolutionary endpoint is a black hole
1.0 × 10^{32}  kilograms Mass of the most massive member of Plaskett’s star
1.2 × 10^{32}  kilograms Rough mass at which a star becomes unstable to pulsations
2 × 10^{33}  kilograms Typical interstellar cloud mass
1 × 10^{36}  kilograms Typical mass of a globular cluster
5 × 10^{36}  kilograms Approximate mass of the Milky Way central black hole
1 × 10^{37}  kilograms Rough baryonic Jeans mass immediately after decoupling
4 × 10^{39}  kilograms Rough mass of the SMC
2 × 10^{40}  kilograms Rough mass of the LMC
2.6 × 10^{42}  kilograms Rough mass of the visible and dark matter in the Milky Way (ApJ 345, 759)
1.3 × 10^{44}  kilograms Rough mass of the stars in the Coma galaxy cluster
6.4 × 10^{44}  kilograms Rough mass of the X-ray gas in the Coma galaxy cluster
2.7 × 10^{45}  kilograms Rough virial mass of the Coma galaxy cluster
6 × 10^{45}  kilograms Rough virial mass of the Abell 2163 galaxy cluster
2 × 10^{48}  kilograms Rough baryonic Jeans mass immediately before decoupling
1.4 × 10^{49}  kilograms Rough total mass in spiral galaxies
2 × 10^{49}  kilograms Rough total (baryonic and nonbaryonic) mass in galaxy clusters
3 × 10^{49}  kilograms Rough total mass in elliptical and spheroidal galaxies
8 × 10^{49}  kilograms Rough total mass of visible matter in the universe
1 × 10^{51}  kilograms Rough total baryon mass predicted by primordial nucleosynthesis
2 × 10^{52}  kilograms Rough total mass of a critical density universe

2 × 10^{−38}  kg m^{−3} Effective density of the 100-300 MHz radio background
1 × 10^{−35}  kg m^{−3} Effective density of the 1–10 MeV γ-ray background
8 × 10^{−35}  kg m^{−3} Effective density of the 2–100 keV X-ray background
1.1 × 10^{−33} kg m^{−3} Upper limit to the effective density of the gravitational wave background
1 × 10^{−32}  kg m^{−3} Effective density of the starlight released in a Hubble time
4.6 × 10^{−31} kg m^{−3} Effective density of the cosmic microwave background radiation
2 × 10^{−29}  kg m^{−3} Smoothed density of visible galactic material throughout universe
2 × 10^{−28}  kg m^{−3} Smoothed baryon density predicted by primordial nucleosynthesis
4.7 × 10^{−27} kg m^{−3} Critical density of the universe = \frac{3H_{0}^{2}}{8πG}
2 × 10^{−24}  kg m^{−3} Typical gas in a cluster of galaxies
3 × 10^{−21}  kg m^{−3} Typical gas in the interstellar medium of the Milky Way
7 × 10^{−21}  kg m^{−3} Dynamically inferred Milky Way disk density
5 × 10^{−20}  kg m^{−3} Typical density of the gas in the central kiloparsec of an interacting or starburst galaxy
1 × 10^{−9}  kg m^{−3} Best room temperature vacuum achieved on Earth
1.7 × 10^{−4} kg m^{−3} Mean density of Antares (19 M_{⊙})
1.3  kg m^{−3} Density of air
700  kg m^{−3} Mean density of Saturn
1 000  kg m^{−3} Density of water
1 300  kg m^{−3} Mean density of Jupiter
1 400  kg m^{−3} Mean density of the Sun
3 300  kg m^{−3} Mean density of the Moon
5 500  kg m^{−3} Mean density of the Earth
7 860  kg m^{−3} Density of iron
19 300  kg m^{−3} Density of gold
5 × 10^{7}  kg m^{−3} Typical white dwarf mean density
3 × 10^{10}  kg m^{−3} Typical white dwarf central density
1.1 × 10^{12}  kg m^{−3} Inverse β decay threshold
4.3 × 10^{14}  kg m^{−3} Neutron drip density
6 × 10^{17}  kg m^{−3} Nuclear density
1 × 10^{18}  kg m^{−3} Typical neutron star central density
5 × 10^{96}  kg m^{−3} Planck-Wheeler density, at which quantum gravitational effects become important = \frac{c^{5}}{G^{2}h̄}
6 × 10^{−25}  joules Lamb-Retherford shift
9.5 × 10^{−25}  joules 21.049 cm photon—hyperfine shift
4 × 10^{−22}  joules Molecular rotation transition
4 × 10^{−21}  joules kT_{room} = translational kinetic energy of atmosphere gas molecule
7 × 10^{−21}  joules Donor level/conduction band gap in a doped semiconductor
3 × 10^{−20}  joules Molecular vibration transition
1.6 × 10^{−19}  joules Valence band/conduction band gap in a semiconductor
2.2 × 10^{−19}  joules Bond energy of an organic substance with half-life 30 000 years
3.8 × 10^{−19}  joules Green light photon
9 × 10^{−19}  joules Valence band/conduction band gap in an insulator
1.1 × 10^{−18}  joules Fermi energy in copper = depth of Fermi sea
1.1 × 10^{−18}  joules Positronium ionization energy
2.2 × 10^{−18}  joules Hydrogen n = 1 binding energy
1.6 × 10^{−16}  joules 1 keV X-ray
1.5 × 10^{−15}  joules Hydrogenic iron n = 1 binding energy
8.18 × 10^{−14}  joules Electron rest mass
1.6 × 10^{−13}  joules 1 MeV γ-ray
1.6 × 10^{−13} (Z_{1} × Z_{2}) joules Coulomb barrier height
1.3 × 10^{−12}  joules Nucleon binding energy
4.3 × 10^{−12}  joules Energy from 4 × (^{1}H) → ^{4}He
1.50 × 10^{−10}  joules Proton rest mass

1.6 × 10^{−7}  joules Particle kinetic energy in a 1 TeV accelerator
10  joules Well hit tennis ball
51  joules Fly’s Eye most energetic cosmic ray event
2 × 10^{5}  joules Energy from a light bulb burning for 1 hour
2 × 10^{7}  joules Rough energy from a 1 kilogram meal
4.2 × 10^{9}  joules Explosion energy of 1 ton of TNT
2 × 10^{10}  joules Rough energy of a lightning flash
2 × 10^{11}  joules Rough human total energy output in a lifetime
1.5 × 10^{14}  joules Typical atomic bomb explosion energy
6 × 10^{16}  joules Tunguska 50 m diameter meteorite impact energy
1 × 10^{17}  joules Powerful H-bomb explosion energy
3 × 10^{17}  joules Elastic wave energy release from a large (M = 8.5) earthquake
5 × 10^{17}  joules Explosion energy of Krakatoa
3 × 10^{18}  joules Superman flying at 70 per cent of light speed
9 × 10^{18}  joules USA electricity usage in 1986
5 × 10^{19}  joules Rough explosion energy of the Lake Toba eruption in Sumatra
2.5 × 10^{22}  joules Energy from the comet Shoemaker-Levy 9 fragment G impact on Jupiter
4 × 10^{23}  joules K-T 10 km diameter meteorite impact energy
1 × 10^{24}  joules Power released by an evaporating black hole during the last second of its life
2.1 × 10^{29}  joules The Earth’s rotational energy
3 × 10^{31}  joules The Earth’s total heat content
7 × 10^{34}  joules Total rotational energy of the planets
2.5 × 10^{35}  joules Rotational energy of the Sun
3 × 10^{36}  joules Gravitational internal binding energy of Jupiter
6 × 10^{37}  joules Nova Persei outburst
3 × 10^{41}  joules A 450 km s^{−1} neutron star kick
3 × 10^{43}  joules Energy needed to make the local bubble
1.3 × 10^{44}  joules Total radiant energy from the Sun — (\frac{1}{10})(0.007)M_{⊙}c^{2}
3 × 10^{44}  joules Energy in photons from a type II supernova explosion
1 × 10^{45}  joules Rough total energy from a cosmological γ-ray burst
3 × 10^{46}  joules Energy in neutrinos from a type II supernova explosion
2 × 10^{53}  joules Typical gravitational binding energy of a galaxy
2 × 10^{54}  joules Typical magnetic and kinetic energy in a large radio lobe (Shu 311)
5.2 × 10^{54}  joules Rotational energy of a 10^{8} M_{⊙} maximal Kerr black hole = 0.29M_{•}c^{2}
5 × 10^{57}  joules Typical gravitational binding energy of a cluster of galaxies
2 × 10^{7}  J kg^{−1} Efficiency of the metabolism of food
1.3 × 10^{13} J kg^{−1} Efficiency of dropping matter onto a white dwarf
6.3 × 10^{14} J kg^{−1} Efficiency of fusion burning hydrogen
5.1 × 10^{15} J kg^{−1} Slowly spiraling accretion onto a Schwarzschild black hole
7.5 × 10^{15} J kg^{−1} Shakura-Sunyaev efficiency for accretion onto a Newtonian “black hole”
8 × 10^{15}  J kg^{−1} Efficiency of dropping matter onto a neutron star
2.6 × 10^{16} J kg^{−1} Slowly spiraling prograde accretion onto a realistically maximally rotating Kerr black hole
3.8 × 10^{16} J kg^{−1} Slowly spiraling prograde accretion onto a theoretically maximally rotating Kerr black hole
8.99 × 10^{16} J kg^{−1} Efficiency of matter-antimatter annihilation
4 × 10^{−6} watts Average energy output from 1 kilogram of the Milky Way
1 × 10^{−3} watts Optical disc player laser
6  watts Amateur short wave radio transmitter
60  watts Light bulb
100  watts Gravitational wave power from the Earth-Sun system
150  watts Human being under normal conditions
750  watts Maximum long duration horse output
1 500  watts Typical fireplace fire

20 000  watts Car
7 × 10^{4}  watts Soft X-ray scattered lunar luminosity
1 × 10^{5}  watts Running Tyrannosaurus Rex
3 × 10^{8}  watts Nuclear power reactor
3 × 10^{8}  watts Moderate thunderstorm electrical power generation rate
1.3 × 10^{9} watts Hoover dam
4 × 10^{9}  watts Soft X-ray Jovian auroral luminosity
3 × 10^{11}  watts USA average electricity usage rate in 1986
1 × 10^{13}  watts Solar radio luminosity
8 × 10^{13}  watts Powerful nanosecond pulse laser
1.7 × 10^{17} watts Insolation of Earth
1 × 10^{20}  watts X-ray luminosity of the quiet Sun
1 × 10^{23}  watts Typical white dwarf luminosity
8 × 10^{24}  watts Gravitational wave radiation from PSR 1913+16
3.9 × 10^{26} watts Solar luminosity
5 × 10^{26}  watts The 2–10 keV flux of the cataclysmic variable FO Aquarii
1 × 10^{28}  watts Typical X-ray luminosity of an X-ray pulsar
1 × 10^{28}  watts X-ray luminosity associated with Sgr A^{∗}
3.5 × 10^{28} watts Rotational energy loss rate from the Geminga pulsar
4.3 × 10^{28} watts Megamaser in NGC 4258
2.4 × 10^{29} watts 5 M_{⊙} star on main sequence
4 × 10^{30}  watts Cygnus X-1 X-ray luminosity
5.5 × 10^{30} watts Luminosity of the 20 M_{⊙} red supergiant Betelgeuse at maximum light
1 × 10^{31}  watts Crab Nebula energy output
2 × 10^{31}  watts Eddington limit for a 1.4 M_{⊙} neutron star
3 × 10^{31}  watts Typical luminosity of Cir X-1 in zero phase
3 × 10^{31}  watts Typical luminosity of GRS 1915+105 in outburst (Nature 371, 46)
1 × 10^{33}  watts Rough luminosity of Eta Carinae in April 1843
5 × 10^{35}  watts Type II supernova peak photon luminosity
3 × 10^{36}  watts Milky Way
1.5 × 10^{38} watts Coma cluster X-ray gas luminosity
1 × 10^{39}  watts 10^{8} M_{⊙} black hole accreting at \frac{1}{10} of the Eddington limit
1 × 10^{39}  watts Typical quasar luminosity
1 × 10^{40}  watts Rough luminosity of the quasar 3C273
2 × 10^{41}  watts Rough luminosity of the z = 2.286 ultraluminous IRAS galaxy IRAS F10214+4724
6 × 10^{43}  watts Rough luminosity of a cosmological γ-ray burst
2 × 10^{46}  watts Rough luminosity of all the stars in the universe
1 × 10^{48}  watts Core collapse neutrino luminosity of a type II supernova (QJRAS 30, 423)
7 × 10^{−7}  kelvins Laser cooling of cesium atoms
1.3 × 10^{−5} kelvins Cosmic microwave background quadrupole anisotropy
5 × 10^{−4}  kelvins Typical Sunyaev-Zel’dovich cosmic microwave background decrement
3.3 × 10^{−3} kelvins Cosmic microwave background dipole anisotropy
0.01  kelvins Typical limit of liquid helium dilution cooling
0.3  kelvins Typical limit of liquid helium evaporation cooling
2.17  kelvins Liquid ^{4}He superfluid transition temperature
2.726  kelvins Cosmic microwave background temperature today
3.20  kelvins Liquid ^{3}He boiling point
4  kelvins Typical limit of Joule-Thomson effect cooling
4.18  kelvins Liquid ^{4}He boiling point
6  kelvins Typical noise temperature of a HEMT receiver at 30 GHz
12  kelvins Lanthanum (under pressure) superconductivity critical temperature (highest for a pure element)
20  kelvins Liquid H_{2} boiling temperature
77  kelvins Liquid N_{2} boiling temperature

133  kelvins Mercury-barium-calcium-copper oxide compound superconductivity critical temperature
273  kelvins Water freezing temperature
311  kelvins Human surface temperature
373  kelvins Water boiling temperature
388  kelvins Brimstone melting temperature—upper limit to the temperature of Hell
506  kelvins Paper burning temperature
740  kelvins Typical surface temperature of Venus
1811  kelvins Melting temperature of iron
3000  kelvins Cosmic microwave background temperature at decoupling
5770  kelvins Solar effective temperature
5 × 10^{5}  kelvins Surface temperature of the Geminga pulsar
3 × 10^{6}  kelvins Polar cap temperature of the Geminga pulsar
3 × 10^{6}  kelvins Typical fusion experiment
1.4 × 10^{7}  kelvins Center of the Sun
1.5 × 10^{7}  kelvins Changeover temperature from the proton-proton chain to the CNO cycle
2.7 × 10^{7}  kelvins Center of a 5 M_{⊙} star
5 × 10^{7}  kelvins Typical gas temperature in a cluster of galaxies
1 × 10^{8}  kelvins Typical cataclysmic variable accretion column shock temperature
4 × 10^{8}  kelvins Characteristic temperature for electron-positron pair production
4 × 10^{8}  kelvins Minimum primordial nucleosynthesis temperature
5 × 10^{8}  kelvins Inner accretion disc temperature of Cyg X-1
7 × 10^{8}  kelvins Thermal electrons become relativistic (v_{e} = \frac{c}{2})
1 × 10^{9}  kelvins Maximum primordial nucleosynthesis temperature
1 × 10^{9}  kelvins Rough superconductivity critical temperature in a neutron star
1 × 10^{10}  kelvins Rough plasma pair catastrophe temperature
3 × 10^{10}  kelvins Core collapse temperature of a supernova (QJRAS 30, 424)
3 × 10^{15}  kelvins Rough electroweak unification temperature
7 × 10^{−13} N m^{−2} Typical gas pressure in a cluster of galaxies
5 × 10^{−12} N m^{−2} Pressure in the best vacuum achieved on Earth
5 × 10^{−6} N m^{−2} Solar radiation pressure at the Earth
1 × 10^{−5} N m^{−2} Pressure of a sound wave at the human threshold of hearing
0.2  N m^{−2} Solar radiation pressure at the surface of the Sun
30  N m^{−2} Pressure of a sound wave at the human threshold of pain
1 × 10^{4}  N m^{−2} Mean human arterial blood overpressure
1.5 × 10^{4} N m^{−2} Standing person
1 × 10^{5}  N m^{−2} Typical atmospheric pressure
1.5 × 10^{6} N m^{−2} High pressure bicycle tire
2 × 10^{7}  N m^{−2} Typical scuba tank pressure
9 × 10^{7}  N m^{−2} Peak pressure of a fist on concrete during a karate strike
1.1 × 10^{8} N m^{−2} Pressure at the bottom of the Marianas trench
6 × 10^{9}  N m^{−2} Pressure needed for the natural crystalization of diamonds
1 × 10^{10}  N m^{−2} Conventional high pressure laboratory press
2 × 10^{11}  N m^{−2} Peak pressure from a diamond anvil pressure cell
5 × 10^{11}  N m^{−2} Central pressure of the Earth
4 × 10^{12}  N m^{−2} Central pressure of Jupiter
2 × 10^{13}  N m^{−2} Radiation pressure at the center of the Sun
2.7 × 10^{16} N m^{−2} Central pressure of the Sun
1 × 10^{23}  N m^{−2} Typical central pressure of a white dwarf
1 × 10^{34}  N m^{−2} Typical central pressure of a neutron star
1 × 10^{−12} W m^{−2} Threshold of hearing (0 dB)
1 × 10^{−11} W m^{−2} Rustle of leaves (10 dB)
1 × 10^{−10} W m^{−2} Quiet whisper (20 dB)

1 × 10^{−9} W m^{−2} Soft music (30 dB)
3 × 10^{−8} W m^{−2} Average residence (45 dB)
1 × 10^{−6} W m^{−2} Background music (60 dB)
3 × 10^{−6} W m^{−2} Ordinary conversation at 0.5 meters (65 dB)
1 × 10^{−5} W m^{−2} Busy street traffic (70 dB)
3 × 10^{−5} W m^{−2} Typical factory (75 dB)
1 × 10^{−3} W m^{−2} Rough boundary of unsafe sound levels (90 dB)
1 × 10^{−2} W m^{−2} Lawn mower (100 dB)
1  W m^{−2} Loud rock music concert (120 dB)
10  W m^{−2} Threshold of pain (130 dB)
100  W m^{−2} Jet plane at 30 meters (140 dB)
9.128 × 10^{−35} kg m^{2} s^{−1} Magnitude of the electron spin angular momentum = \sqrt{\frac{3}{4}}h̄
1.054 × 10^{−34} kg m^{2} s^{−1} Planck angular momentum = h̄
3 × 10^{12}  kg m^{2} s^{−1} Rough angular momentum of an artificial satellite
5.9 × 10^{33}  kg m^{2} s^{−1} Earth rotational angular momentum
1.7 × 10^{41}  kg m^{2} s^{−1} Sun rotational angular momentum
3.1 × 10^{43}  kg m^{2} s^{−1} Total angular momentum of our solar system
2 × 10^{55}  kg m^{2} s^{−1} Angular momentum of the Sun around the center of the Milky Way
4 × 10^{66}  kg m^{2} s^{−1} Rough total angular momentum of the Milky Way
1 × 10^{−13} teslas Spontaneous human brain activity
1 × 10^{−12} teslas Evoked human brain activity
1 × 10^{−12} teslas Typical magnetic field needed for good radio reception
1 × 10^{−11} teslas Typical magnetic field from a human heart
1 × 10^{−10} teslas Typical 50/60 Hz magnetic field inside a building
5 × 10^{−10} teslas Typical magnetic field strength in the local interstellar medium
1 × 10^{−9} teslas Typical magnetic field strength in a radio lobe
1 × 10^{−8} teslas Typical magnetic field strength in the central 300 parsecs of the Milky Way
5 × 10^{−8} teslas Magnetic field in the Crab Nebula
1 × 10^{−7} teslas Magnetic field in The Arc at the Milky Way center
1 × 10^{−6} teslas Typical magnetic field from a hand held cordless phone
3 × 10^{−5} teslas Magnetic field at Earth’s surface
1 × 10^{−4} teslas Magnetic field near Sun’s pole
4 × 10^{−4} teslas Magnetic field at Jupiter’s cloud tops
0.1  teslas Ap star magnetic field
0.2  teslas Sunspot magnetic field
1  teslas Typical medical NMR magnetic field
2  teslas Magnetic field felt by the electron in an n = 1 hydrogen atom
2  teslas RS CVn star spot magnetic field
12  teslas Typical magnetic field used in high resolution NMR spectroscopy
25  teslas Powerful superconducting/normal hybrid magnet
1 × 10^{8}  teslas Typical single pulsar dipole magnetic field strength
4.4 × 10^{9} teslas Magnetic field strength when the energy of the first electron Landau level is

comparable to the electron rest mass (MNRAS 275, 257)

1 × 10^{4} amperes Peak lightning bolt current
5 × 10^{6} amperes Current along Io’s flux tube (Science 262, 1035)
3 × 10^{8} volts Typical potential difference across a lightning bolt
3 × 10^{6} V m^{−1} Electrical discharge field in air with ions
1 × 10^{12} V m^{−1} Typical electric field at the surface of a pulsar
1 × 10^{19} V m^{−1} QED pair production electric field limit

240  bits Eigen replicator genome — 120 nucleotides
440  bits Spiegelman monster genome — 220 nucleotides
9 000  bits Q_{β} virus genome — 4 500 nucleotides
32 000  bits Single spaced typed page
6.6 × 10^{5} bits Uncompressed ASCII version of Lysistrata
4.36 × 10^{6} bits C. elegans nematode chromosome III — longest contiguous piece of DNA sequenced
6 × 10^{6}  bits E. Coli genome — 3 million nucleotides
6.3 × 10^{6} bits Uncompressed ASCII version of A Tale of Two Cities
2.6 × 10^{7} bits Yeast genome — 13 million nucleotides
3.6 × 10^{7} bits Uncompressed ASCII version of The Bible
2 × 10^{8}  bits Nematode genome — 100 million nucleotides
1 × 10^{9}  bits Microcomputer hard drive
2.5 × 10^{9} bits Twenty volume edition of the Oxford English Dictionary on CD-ROM
7 × 10^{9}  bits MACHO project data intake per night
9 × 10^{9}  bits Human genome
8 × 10^{10}  bits Sloan Digital Sky Survey spectral database
1.4 × 10^{11} bits Cambridge X-ray group total disk space
2 × 10^{11}  bits International Halley watch database
5 × 10^{11}  bits VCR tape
1 × 10^{12}  bits Speed reading for 80 years with perfect memory
8.8 × 10^{12} bits HST archive size in mid-1994
9 × 10^{12}  bits VLA FIRST Survey data set
2 × 10^{13}  bits Digitized new Palomar sky survey
5 × 10^{13}  bits Sloan Digital Sky Survey 0.4 arcsecond pixel map
8 × 10^{13}  bits NSF backbone traffic in March 1994
3 × 10^{16}  bits HDTV for 80 years with perfect memory
200  bits s^{−1} Typical reading and talking rate
500  bits s^{−1} Fast speed reading rate
24 000  bits s^{−1} Hipparcos satellite data acquisition rate
56 000  bits s^{−1} Typical Internet link
1 × 10^{5}  bits s^{−1} High quality audio
1.4 × 10^{5} bits s^{−1} HST archive mean growth rate
1 × 10^{7}  bits s^{−1} Xerox Ethernet
1 × 10^{7}  bits s^{−1} HDTV bit rate after compression
4.5 × 10^{7} bits s^{−1} Internet backbone T3 link
6 × 10^{7}  bits s^{−1} Sloan Digital Sky Survey data acquisition rate
3 × 10^{10} bits s^{−1} Soliton optical fiber information transmission rate
2 × 10^{15} bits s^{−1} Sexual reproduction genetic information transmission rate
2 100  1994 USA dollars Average USA public school spending per pupil in 1960
6 200  1994 USA dollars Average USA public school spending per pupil in 1993
2 × 10^{5}  1994 USA dollars Vancouver liquid mercury mirror telescope
9 × 10^{6}  1994 USA dollars Income of St. John’s College, Cambridge during 1992–1993
1.5 × 10^{7} 1994 USA dollars Development and construction cost of the Giant Meterwave Radio Telescope near

Pune, India

2.9 × 10^{7} 1994 USA dollars Total predicted cost of the Sloan Digital Sky Survey
3 × 10^{7}  1994 USA dollars Rough United States ACLU yearly national operations budget
3 × 10^{7}  1994 USA dollars Rough United States NRA lobbying budget
5.5 × 10^{7} 1994 USA dollars Development, construction, and launch cost of the Clementine I spacecraft
8.5 × 10^{7} 1994 USA dollars VLBA development and construction cost
9 × 10^{7}  1994 USA dollars Development and construction cost of the Keck telescope
1 × 10^{8}  1994 USA dollars Development and construction cost of the ASCA X-ray satellite
1.5 × 10^{8} 1994 USA dollars Typical holdings of an old wealthy Oxbridge college

1.5 × 10^{8} 1994 USA dollars Rough cost of a European Ariane rocket launch
2 × 10^{8}  1994 USA dollars Development and construction cost of the VLA
2.1 × 10^{8} 1994 USA dollars Total spending in the 1994 United States senate election campaigns
3.7 × 10^{8} 1994 USA dollars Rough construction and operation cost (for 4 years) of LIGO
7.8 × 10^{8} 1994 USA dollars Rough United States NEA revenues
9 × 10^{8}  1994 USA dollars Development, construction, and launch cost of the Magellan probe
9.8 × 10^{8} 1994 USA dollars Development, construction, and launch cost of the Mars Observer spacecraft
1.1 × 10^{9} 1994 USA dollars Worldwide Visa and MasterCard fraud in 1993
1.5 × 10^{9} 1994 USA dollars Planned total cost of the AXAF X-ray mission
1.6 × 10^{9} 1994 USA dollars Estimated cost of the CERN Large Hadron Collider
1.7 × 10^{9} 1994 USA dollars Barclays 1994 half-year profits
1.8 × 10^{9} 1994 USA dollars Amount of food stamp fraud in the USA in 1993
3.5 × 10^{9} 1994 USA dollars Planned total cost of the Cassini spacecraft
3.8 × 10^{9} 1994 USA dollars Microsoft revenue in 1993
1 × 10^{10}  1994 USA dollars Rough monetary losses associated with BCCI
1.3 × 10^{10} 1994 USA dollars Lockheed revenue in 1993
1.4 × 10^{10} 1994 USA dollars NASA planned 1995 budget
1.5 × 10^{10} 1994 USA dollars Rough United Nations yearly budget
2.8 × 10^{10} 1994 USA dollars Planned cost for the space station
4.4 × 10^{10} 1994 USA dollars Total cost of the first 20 B-2 Stealth bombers
5 × 10^{10}  1994 USA dollars Clinton administration 1995 loan guarantee to Mexico
1 × 10^{11}  1994 USA dollars Rough annual gap between USA IRS estimates of federal taxes due and tax revenues

actually collected

2.6 × 10^{11} 1994 USA dollars United States 1994 military spending
2.6 × 10^{11} 1994 USA dollars United States 1994 predicted defecit
8 × 10^{11}  1994 USA dollars United States 1994 entitlement spending (e.g. social security, medicare, medicaid,

food stamps, unemployment compensation, farm aid, federal pensions)

1 × 10^{12}  1994 USA dollars Rough total United States health care spending in 1994
1.3 × 10^{12} 1994 USA dollars United States 1994 tax receipts
1.5 × 10^{12} 1994 USA dollars United States 1994 federal government spending
3.5 × 10^{12} 1994 USA dollars United States antipoverty spending since 1965
4.4 × 10^{12} 1994 USA dollars United States 1994 national debt
6.4 × 10^{12} 1994 USA dollars United States 1994 gross domestic product
1.4 × 10^{13} 1994 USA dollars United States 1994 unfunded liabilities for social security, medicare, federal

civil-service retirement and military retirement

3.1 × 10^{14} 1994 USA dollars World 1994 GDP
0.004  Black velvet optical light albedo
0.012  Black cloth optical light albedo
0.04  Halley’s comet optical light albedo
0.068  Moon optical light albedo
0.14  Mars optical light albedo
0.40  Earth optical light albedo
0.80  White paper optical light albedo
0.85  Venus optical light albedo
1 × 10^{−10} pc^{−3} Space density of X-ray binaries in the Milky Way
5 × 10^{−7} pc^{−3} Space density of pulsars in the solar neighborhood
3 × 10^{−6} pc^{−3} Space density of cataclysmic variables in the solar neighborhood
0.1  pc^{−3} Space density of main sequence stars in the solar neighborhood
1 × 10^{−9} Mpc^{−3} Ultraluminous IRAS galaxies
1 × 10^{−9} Mpc^{−3} Quasars
1 × 10^{−7} Mpc^{−3} QSOs

1 × 10^{−7} Mpc^{−3} FR II radio galaxies (edge-brightened)
5 × 10^{−6} Mpc^{−3} FR I radio galaxies (not edge-brightened)
1 × 10^{−4} Mpc^{−3} Seyfert galaxies
0.1  Mpc^{−3} Field galaxies
0.6  Mpc^{−3} Group galaxies
−26.8  V apparent magnitude of the Sun
−12.5  V apparent magnitude of the full Moon
−4.4  V apparent magnitude of Venus at its brightest
−1.5  V apparent magnitude of Sirius
0.1  V apparent magnitude of the LMC
0.5  V apparent magnitude of Betelgeuse
3.4  V apparent magnitude of M31
6.5  V apparent magnitude of naked eye limit at a dark site
9  V apparent magnitude of 50-mm binocular limit
10.5  V apparent magnitude of the z = 0.00363 Seyfert 2 NGC 1068
11.6  V apparent magnitude of the z = 0.00170 H II region galaxy NGC 5408
12.8  V apparent magnitude of the z = 0.158 quasar 3C273
13.6  V apparent magnitude of the z = 0.0167 Seyfert 1 NGC 7469
14.5  V apparent magnitude of 12-in telescope visual limit
15.3  V apparent magnitude of the z = 0.239 quasar PG 0953+414 (ApJ 435, L12)
19.55  V apparent magnitude of the z = 3.67 quasar DHM 0054 − 284
23.5  V apparent magnitude of 200-in telescope photographic limit
25.5  V apparent magnitude of a typical Cepheid in M100 (ApJ 435, L33)
29.5  V apparent magnitude limit of an 18-hour exposure with HST
−27.5  M_{V} of the z = 0.158 quasar 3C273
−24.1  M_{V} of the z = 0.239 quasar PG 0953+414 (ApJ 435, L12)
−21.1  M_{V} of M31
−21.0  M_{V} of a typical Seyfert host galaxy
−19.5  M_{V} limit on the host galaxy of the z = 0.239 quasar PG 0953+414 (ApJ 435, L12)
−18.5  M_{V} of the LMC
−6.0  M_{V} of Betelgeuse
4.4  M_{V} of the most luminous star in Alpha Centauri
4.79  M_{V} of the Sun

Note: I assume an inflationary (Ω_{0} = 1.0) dust-filled Friedmann-Robertson-Walker universe with H_{0} = 50 km s^{−1} Mpc^{−1} = 1.62 × 10^{−18} s^{−1} for all calculations. Space densities are given for “here and now.”